5. A teacher wants to test his assumption that less than 30% of the Senior High


School students liked research class. After randomly collecting 150 samples,


he found out that only 40 students like their research class.


Claim:


Parameter


Symbol for parameter


Ho and Ha complementary pair.


Hypotheses in words:


Ho


Ha:


Hypotheses in symbols


Ho


На:

1
Expert's answer
2022-05-20T06:50:09-0400

The population proportion "p" is the parameter.

Claim "p<0.3."

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:" the population proportion is at least 30%.

"H_a:" the population proportion is less than 30%.

"H_0:p\ge0.3"

"H_a:p<0.3"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c =-1.6469."

The rejection region for this left-tailed test is "R = \{z: z<-1.6469\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p}{\sqrt{\dfrac{p(1-p)}{n}}}=\dfrac{\dfrac{40}{150}-0.3}{\sqrt{\dfrac{0.3(1-0.3)}{150}}}=-0.891"

Since it is observed that "z =-0.891>-1.6469=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z<-0.891)=0.186465," and since "p = 0.186465 > 0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than 0.3, at the "\alpha = 0.05" significance level.